OpenStudy (ggdvaaa):

A drilling machine can drill 12 holes in 6 minutes. If the machine drills holes at a constant speed, it can drill ______ holes in 25 minutes.

OpenStudy (ggdvaaa):

@Directrix

OpenStudy (anonymous):

If the drill drills 12 holes in 6 minutes how many holes can it drill in one minute?

OpenStudy (ggdvaaa):

2 ?

rebeccaxhawaii (rebeccaxhawaii):

brills got it

OpenStudy (anonymous):

Yup because 12 divided by 6 is 2 So if it drills 2 holes a minute and there are 25 minutes how many holes would it drill?

OpenStudy (ggdvaaa):

uh 125?

OpenStudy (ggdvaaa):

he left me .

OpenStudy (anonymous):

Well not quite So if it drills 2 holes in 1 minute and you want how many holes in 25 mins right? Thats just\[\frac{ 2 }{ 1 }=\frac{ x }{ 25 }\]Where x is the number of holes

jhonyy9 (jhonyy9):

use the rule of 3 simple 12 holes --- 6 minutes x holes ---- 25 minutes ---------------------- x = 25*12/6 = ?

OpenStudy (ggdvaaa):

300?

jhonyy9 (jhonyy9):

why 300 ? 12/6=2 and 2*25 = ?

OpenStudy (ggdvaaa):

oh i did 2 times 25 pff

OpenStudy (ggdvaaa):

but you divide right?

OpenStudy (anonymous):

Well look at my equation @ggdvaaa if you simplify it you get\[2=\frac{ x }{ 25 }\] how would you solve that?

OpenStudy (ggdvaaa):

x is 6? im sorry im slow tbh

OpenStudy (anonymous):

How would you get the x by itself?

OpenStudy (ggdvaaa):

im not sure.

OpenStudy (ggdvaaa):

i guessed

OpenStudy (anonymous):

You multiply both sides by 25 right? Because that eliminates the 25 under the x So what would x=?

OpenStudy (ggdvaaa):

both sides of

OpenStudy (anonymous):

both sides of the equation So you would do\[25 \times 2=25 \times \frac{ x }{ 25 }\] What would that be?